# Occlusion-Aware Ground Target Tracking by a Dubins Vehicle Using Visibility Volumes

### v1

@@ 4.2 Environment Model
The UAV operates in an urban environment that consists of a ground plane with a road network and a collection of 2.5-dimensional objects representing buildings or other structures, $B$. The $i$th object, in the set $B$ of $N_{B}$ objects, is an extruded polygon $B_{i}=\{(x,y,z)\in\mathbb{R}^{3}~{}|~{}(x,y)\in A_{i}$ and $z\in[0,h_{i}]\}$, where $A_{i}$ is a simple two-dimensional polygon, and $h_{i}$ is the height of the obstacle. The boundary of polygon $A_{i}$ is denoted $\partial A_{i}$, an ordered set of points with a positive signed area. The set of points on the interior of $A_{i}$ is denoted ${\rm int}(A_{i})$. The polygonal areas of each object do not intersect, ${\rm int}(A_{i})\cap{\rm int}(A_{j})=\emptyset$ for all $i\neq j$ with $i,j\in\{0,\dots,N_{B}-1\}$ where $\emptyset$ is the empty set. The polygons rest on the ground plane, $z=0$ (changes in elevation are not considered). An example 2.5D environment is shown in Fig. 1.
The road network is a graph $G$ with a set of nodes $N$ and bidirectional edges $E$. Each node is assigned a position along the road and edges are assigned between adjacent nodes when connected by a road. This allows the POI to move between nodes on the graph to approximate traveling along a road. The $i$th node has a corresponding point $\bm{p}_{i}\in\mathbb{R}^{2}$ located on the ground plane in the 2.5D urban environment. Each node is unique $\bm{p}_{i}\neq\bm{p}_{j}$. The edges, defined by pairs of nodes $(i,j)$, cannot intersect with any obstacle, $l(\bm{p}_{i},\bm{p}_{j})\cap A_{k}=\emptyset$ for all $(i,j)\in E$ and $k\in\{0,\dots,N_{B}-1\}$ where $l(\bm{p}_{i},\bm{p}_{j})=\{\bm{p}_{i}+\tau(\bm{p}_{j}-\bm{p}_{i})~{}{\rm for}~{}\tau\in[0,1]\}$ is the line segment between nodes. A road network where nodes are only defined at intersections of two roads is shown in Fig. 1.
Figure 1: An example of the environment where a UAV tracks a point of interest. Left: a two-dimensional view of the urban environment with the building footprints (blue) and the roads (magenta). The circles are intersections where the roads meet. Right: a three-dimensional view of the same urban environment. The gray is the ground plane and the buildings are colored yellow to purple to indicate height.
@@ 4.3 Point of Interest Model
for $t_{i}\in[t_{i},t_{j}]$ where $t_{i}$ is the time the POI is at the $i$th node. The POI will travel along this edge until it reaches the $j$th node at time
$t_{j}=\frac{||\bm{p}_{j}-\bm{p}_{i}||}{v_{g}}+t_{i}\;.$
When the POI reaches the $j$th node, the POI transitions to a new edge $(j,k)\in E$ such that $k\neq j$.
@@ 4.4 Measurement Model
The UAV is equipped with an imaging sensor (e.g., RGB, infrared, or hyperspectral camera) mounted on a three-axis pan-tilt-roll gimbal that can regulate the camera to point at a position on the ground plane. The camera can measure the POI when it is within sensing range and not obscured by obstacles. The region of space where the camera may image a point $\bm{g}$ on the ground plane is called a visibility volume (VV). When the POI is moving, the VV is time-varying. To approximate the time-varying VV, a discrete set of VVs is created where each VV corresponds to a different time. The line segment between the UAV at position ${\bm{\rho}}$ and the POI at position $\bm{g}$ is
$L(\tau;{\bm{g}},{\bm{\rho}})=\tau\begin{bmatrix}g_{x}\\ g_{y}\\ 0\end{bmatrix}+(1-\tau)\begin{bmatrix}\rho_{x}\\ \rho_{y}\\ \rho_{z}\end{bmatrix}\;,$
where $\tau\in[0,1]$ is a parameter. The VV is a set of points ${\bm{\rho}}=[\rho_{x},\rho_{y},\rho_{z}]^{\rm T}\in F$ that have a direct line-of-sight to each target (i.e., not obscured by buildings) and satisfy other sensing constraints. For a gimbaled camera, the VV for a target located at ${\bm{g}}$ is the subset of the feasible airspace $F$ lies within direct line-of-sight to the target, and within a maximum range $d_{\rm max}$ relative to the target:
@@
For brevity, visibility volumes (5) are henceforth denoted $\alpha({\bm{g}})$. The maximum range $d_{\rm max}$ is determined by the minimum pixel requirements for imaging a target of a known size with a particular camera resolution. VVs are computed using [6], resulting in a triangular mesh.
@@ 4.5 Problem Statement: Reference Orbit Design
An orbit may serve as a reference path for a UAV continuously encircling the POI. The reference orbit should be inside the POI’s VV as the POI travels through the environment over a time interval $[t_{0},t_{f}]$ where $t_{f}>t_{0}$. By following the VO, inside the VV, the UAV maintains a view of the target at all times. Let $O(\theta,t)~{}:~{}\mathbb{S}\times\mathbb{R}\to\mathbb{R}^{2}$ denote a time-varying orbit that is parameterized by the angle $\theta$ about the moving point $\bm{g}(t)=[g_{x}(t)~{}g_{y}(t)]^{\rm T}$ and the time $t$. For each point in time $t$, the orbit function $O(\theta,t)$ produces a closed non-self-intersecting curve in the $xy$ plane when sweeping out the angle $\theta\in\mathbb{S}$. The orbit is centered on ${\bm{g}}(t)$ and is located at a constant altitude $z=h_{\rm UAV}$. The curvature of the orbit in the inertial frame, denoted $\kappa(O(\theta,t),\theta,t)$ at any angle $\theta$ and time $t$, must remain less than the maximum curvature $\kappa_{\rm max}$ to be feasible for the system (1). The general form of the curvature for a 2D curve is given in (4). The orbit must also remain in the POI’s visibility volume, $\alpha(\bm{g}(t))$. While many types of orbit geometries are feasible, this work considers circular orbits. The time-varying orbit is defined as
@@
for $\theta\in\mathbb{S}$ where $R(t)$ is a time-varying radius. The standoff distance (i.e., the radius of the orbit) is maximized to reduce the curvature and observe the POI from a shallower angle at a greater distance. The optimization problem is to specify $R(t)$ so that the orbit remains feasible for the Dubins vehicle model (1), guarantees the visibility of the POI, and maximizes the standoff distance:
@@
where $\bm{g}(t)$ is a known continuous path through the environment, $\bm{p}_{i}$ is the location of the $i$th (initial) node on the road network, and $\bm{p}_{j}$ is the final node. The first constraint (8) ensures the orbit is contained inside the visibility volume. The curvature constraint (9) ensures that the UAV’s turn rate constraint (2) is satisfied on the orbit. The final constraints (10) and (11) specify the POI start and end positions on the road network $G$.
@@ 4.6 Problem Statement: Orbit Controller Design
The control objective is to design a controller $u_{\psi}(t,\bm{q})$ that steers the system (1) onto the orbit $O(\theta,t)$,
$\lim_{t\rightarrow\infty}~{}[q_{x}(t),q_{y}(t)]^{\rm T}-O_{\rm min}(t,\bm{q}(t))\rightarrow\bm{0}\;,$
where $O_{\rm min}(t,\bm{q}(t))\in\mathbb{R}^{2}$ is the closest point on the orbit to the configuration $\bm{q}(t)$, and the point $[q_{x}(t),q_{y}(t)]^{\rm T}$ is the inertial position component of the configuration $\bm{q}(t)$. In the case where the UAV is in the middle of the circular orbit, multiple points are equidistant from the internal position $[q_{x}(t),q_{y}(t)]^{\rm T}$ and the UAV is attracted to the point on the orbit in the direction of the UAV’s heading.
@@ 5 Moving View Polygons
First, this section shows the conditions under which a given orbit $R(t)$ (6) is feasible for the Dubins vehicle (1). Then, to better optimize the orbit $R(t)$, an adaptive discretization approach determines the locations $\mathcal{P}$ at which to compute the VVs by avoiding drastic changes in VV geometry. The VVs are then sliced along a constant altitude plane $z=h_{\rm UAV}$ with [56] to produce the VPs. Finally, an algorithm is presented that creates the final time-varying orbit along the POI’s path while considering feasibility constraints. From this section forward, function parameters will be suppressed for brevity.
@@ 5.1 Feasibility of Reference Orbits
$\dot{\theta}=\frac{-\dot{\bm{g}}\cdot\bm{e}_{\theta}\pm\sqrt{v^{2}-(\dot{\bm{g}}\cdot\bm{e}_{r}+\dot{R})^{2}}}{R}\;,$ (14)
where the positive solution represents counter-clockwise (CCW) motion about $\bm{g}$ and the negative solution represents clockwise (CW) motion.
@@
$\dot{\theta}=\frac{-\dot{\bm{g}}\cdot\bm{e}_{\theta}\pm\sqrt{(\dot{\bm{g}}\cdot\bm{e}_{r})^{2}-\left(v_{g}^{2}-v^{2}+2\dot{R}\dot{\bm{g}}\cdot\bm{e}_{r}+\dot{R}^{2}\right)}}{R}\;.$ (20)
@@
Equation (14) is a time-varying differential equation valid over the interval $t\in[t_{i},t_{j}]$ for any initial condition $\theta(t_{i})=\theta_{i}$ and direction (clockwise or counter-clockwise) that defines a unique constant-speed trajectory remaining on the locus of orbits defined by $R$ and $\bm{g}$. The vehicle heading on this orbit is implicitly determined by the resulting path.
@@
Let “remain on the orbit $O$” be the condition wherein the angular velocity $\dot{\theta}$ about $\bm{g}$ described by (14) is greater or equal to zero for counter-clockwise circulation and less than or equal to zero for clockwise circulation for all $\theta\in\mathbb{S}$, and $t\in[t_{i},t_{j}]$.
@@
$\ddot{\bm{x}}_{\mathcal{I}}$ $=\ddot{x}_{\mathcal{I},r}\bm{e}_{r}+\ddot{x}_{\mathcal{I},\theta}\bm{e}_{\theta}$
$=(\dot{a}-b\dot{\theta})\bm{e}_{r}+(\dot{b}+a\dot{\theta})\bm{e}_{\theta}\;,$ (23)
@@
A similar transformations relates the velocities $\dot{x}_{\mathcal{I},r}$ and $\dot{x}_{\mathcal{I},\theta}$ with $\dot{x}_{\mathcal{I},x}$ and $\dot{x}_{\mathcal{I},y}$. Therefore,
$\kappa_{s}$ $=\frac{\ddot{x}_{\mathcal{I},y}\dot{x}_{\mathcal{I},x}-\ddot{x}_{\mathcal{I},x}\dot{x}_{\mathcal{I},y}}{\left(\dot{x}_{\mathcal{I},x}^{2}+\dot{x}_{\mathcal{I},y}^{2}\right)^{3/2}}$
$=\Big{[}(\ddot{x}_{\mathcal{I},r}\sin{\theta}+\ddot{x}_{\mathcal{I},\theta}\cos{\theta})(\dot{x}_{\mathcal{I},r}\cos{\theta}-\dot{x}_{\mathcal{I},\theta}\sin{\theta})-(\ddot{x}_{\mathcal{I},r}\cos{\theta}-\ddot{x}_{\mathcal{I},\theta}\sin{\theta})(\dot{x}_{\mathcal{I},r}\sin{\theta}+\dot{x}_{\mathcal{I},\theta}\cos{\theta})\Big{]}v^{-3/2}$
$=\Big{[}\ddot{x}_{\mathcal{I},r}\dot{x}_{\mathcal{I},r}\sin\theta\cos\theta-\ddot{x}_{\mathcal{I},r}\dot{x}_{\mathcal{I},r}\sin\theta\cos\theta-\ddot{x}_{\mathcal{I},r}\dot{x}_{\mathcal{I},\theta}\sin^{2}\theta+\ddot{x}_{\mathcal{I},\theta}\dot{x}_{\mathcal{I},r}\cos^{2}\theta$
$-\ddot{x}_{\mathcal{I},\theta}\dot{x}_{\mathcal{I},\theta}\sin\theta\cos\theta+\ddot{x}_{\mathcal{I},\theta}\dot{x}_{\mathcal{I},\theta}\sin\theta\cos\theta-\ddot{x}_{\mathcal{I},r}\dot{x}_{\mathcal{I},\theta}\cos^{2}\theta+\ddot{x}_{\mathcal{I},\theta}\dot{x}_{\mathcal{I},r}\sin^{2}\theta\Big{]}\left(\dot{x}_{\mathcal{I},r}^{2}+\dot{x}_{\mathcal{I},\theta}^{2}\right)^{-3/2}$
$=\frac{\ddot{x}_{\mathcal{I},\theta}\dot{x}_{\mathcal{I},r}-\ddot{x}_{\mathcal{I},r}\dot{x}_{\mathcal{I},\theta}}{\left(\dot{x}_{\mathcal{I},r}^{2}+\dot{x}_{\mathcal{I},\theta}^{2}\right)^{3/2}}\;.$ (25)
@@
$=\frac{1}{v^{3}}\left(\frac{-\dot{a}a^{2}}{b}-\frac{\dot{a}b^{2}}{b}\right)+\frac{\dot{\theta}}{v}$
$=-\frac{\dot{a}}{b}\frac{a^{2}+b^{2}}{v^{3}}+\frac{\dot{\theta}}{v}$
@@
If the system (1) remains on the orbit $O$ then the unsigned curvature of the path (3) is given by
@@
Lemma 2 applies to system (1), because system (1) has constant velocity; Therefore, the signed curvature of a trajectory is given by (32) and the unsigned curvature is
@@
Note that the heading history of the system (1) is implicitly defined by a trajectory $\bm{x}_{\mathcal{I}}\in\mathbb{R}^{2}$. If a trajectory $\bm{x}_{\mathcal{I}}$ has constant speed $v$ and bounded unsigned curvature $\kappa\leq\kappa_{\rm max}$ then the system (1) can follow the trajectory. The Dubins model can be written as
$\begin{bmatrix}\dot{q}_{x}\\ \dot{q}_{y}\\ \dot{q}_{\psi}\end{bmatrix}=\begin{bmatrix}v\cos{\psi}\\ v\sin{\psi}\\ v\kappa_{s}\end{bmatrix}\;,$
where $v\kappa_{s}=u_{\psi}$ and $\kappa_{s}$ is the signed curvature. Therefore, a trajectory is uniquely defined by an initial condition,
$\bm{q}=\begin{bmatrix}q_{x,0}\\ q_{y,0}\\ q_{\psi,0}\end{bmatrix}\;,$
and a curvature control $|\kappa_{s}|<\kappa_{\rm max}$.
@@
Let the distance from the orbit be
$\tilde{r}=R-\sqrt{(g_{x}-q_{x})^{2}+(g_{y}-q_{y})^{2}}\;.$
The first derivative of the distance is
@@
$\ddot{\tilde{r}}=\dot{\theta}\dot{\bm{g}}\cdot\bm{e}_{\theta}\left(\frac{v}{\pm\sqrt{v^{2}-\left(\dot{\bm{g}}\cdot\bm{e}_{\theta}+\dot{R}\right)^{2}}}\right)\sin(q_{\psi}-\theta)-\dot{\theta}\dot{\bm{g}}\cdot\bm{e}_{\theta}\;.$ (47)
@@
$\ddot{\tilde{r}}$ $=\dot{\theta}\dot{\bm{g}}\cdot\bm{e}_{\theta}\left(\frac{v}{\pm\sqrt{v^{2}-\left(\dot{\bm{g}}\cdot\bm{e}_{\theta}+\dot{R}\right)^{2}}}\right)\left(\frac{\pm\sqrt{v^{2}-\left(\dot{\bm{g}}\cdot\bm{e}_{\theta}+\dot{R}\right)^{2}}}{v}\right)-\dot{\theta}\dot{\bm{g}}\cdot\bm{e}_{\theta}$
$=0\;.$
@@
Suppose that the system (1) begins on an orbit $O$ with initial conditions (39)-(41). The system (1) remains on the orbit $O$ if and only if the curvature control is given by the left-hand side (LHS) of (32) and the inequality of (32) is satisfied.
@@
Part (ii): From the initial condition ${\bm{q}}_{0}$ the polar angle $\theta_{0}$ is known and the initial position is $O(\theta_{0},t_{0})$. From Lemma 4, if system (1) follows the curvature control (32) with the assumed initial conditions, then system (1) will remain on the orbit. ∎
@@
If orbit $O$ satisfies (32) then $\kappa$ is defined by the LHS of (32). For (32) to be real-valued it is required that
$v^{2}-(\dot{\bm{g}}\cdot\bm{e}_{r}+\dot{R})^{2}\geq 0\;.$
Moreover, this implies that
$-v$ $\leq\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\leq v$
@@
Corollary 1.
As the contrapositive of Lemma 5, if the rate-of-change for the radius $R$ is not bounded by (51) then the curvature of the orbit $O$ is not bounded by (32).

### v2

@@ 3.2 Environment Model
The UAV operates in an urban environment that consists of a ground plane with a road network and a collection of 2.5-dimensional objects representing buildings or other structures, $B$. The $i$th object, in the set $B$ of $N_{B}$ objects, is an extruded polygon $B_{i}=\{(x,y,z)\in\mathbb{R}^{3}~{}|~{}(x,y)\in A_{i}$ and $z\in[0,h_{i}]\}$, where $A_{i}$ is a simple two-dimensional polygon, and $h_{i}$ is the height of the obstacle. The boundary of polygon $A_{i}$ is denoted $\partial A_{i}$, an ordered set of points with a positive signed area. The set of points on the interior of $A_{i}$ is denoted ${\rm int}(A_{i})$. The polygonal areas of each object do not intersect, ${\rm int}(A_{i})\cap{\rm int}(A_{j})=\emptyset$ for all $i\neq j$ with $i,j\in\{0,\dots,N_{B}-1\}$ where $\emptyset$ is the empty set. The polygons rest on the ground plane, $z=0$ (changes in elevation are not considered). An example 2.5D environment is shown in Fig. 1, where the ground plane is gray and the buildings are colored yellow to purple based on height.
The road network is a graph $G$ with a set of nodes $N$ and bidirectional edges $E$. Each node is assigned a position along the road and edges are assigned between adjacent nodes when connected by a road. This allows the POI to move between nodes on the graph to approximate traveling along a road. The $i$th node has a corresponding point $\bm{p}_{i}\in\mathbb{R}^{2}$ located on the ground plane in the 2.5D urban environment. Each node is unique $\bm{p}_{i}\neq\bm{p}_{j}$. The edges, defined by pairs of nodes $(i,j)$, cannot intersect with any obstacle, $l(\bm{p}_{i},\bm{p}_{j})\cap A_{k}=\emptyset$ for all $(i,j)\in E$ and $k\in\{0,\dots,N_{B}-1\}$, where $l(\bm{p}_{i},\bm{p}_{j})=\{\bm{p}_{i}+\tau(\bm{p}_{j}-\bm{p}_{i})~{}{\rm for}~{}\tau\in[0,1]\}$ is the line segment between nodes. A road network where nodes are only defined at intersections of two roads is shown in Fig. 1, where the footprints of the buildings are blue and the road network is magenta.
Figure 1: An example of the environment where a UAV tracks a point of interest.
@@ 3.3 Point of Interest Model
for $t_{i}\in[t_{i},t_{j}]$ where $t_{i}$ is the time the POI is at the $i$th node. The POI will travel along this edge until it reaches the $j$th node. When the POI reaches the $j$th node, the POI transitions to a new edge $(j,k)\in E$ such that $k\neq j$.
@@ 3.4 Measurement Model
The UAV is equipped with an imaging sensor (e.g., RGB, infrared, or hyperspectral camera) mounted on a three-axis pan-tilt-roll gimbal that can regulate the camera to point at the ground plane. The camera can measure the POI when it is within sensing range and not obscured by obstacles. The region of space where the camera may image a point $\bm{g}$ is called a visibility volume (VV). When the POI is moving, the VV is time-varying. To approximate the time-varying VV, a discrete set of VVs is created where each VV corresponds to a different time. The line segment between the UAV at position ${\bm{\rho}}$ and the POI at position $\bm{g}$ is $L(\tau;{\bm{g}},{\bm{\rho}})=\tau[{\bm{g}}^{\rm T}~{}0]^{\rm T}+(1-\tau){\bm{\rho}}\;,$ where $\tau\in[0,1]$ is a parameter. The VV is a set of points ${\bm{\rho}}=[\rho_{x},\rho_{y},\rho_{z}]^{\rm T}\in F$ that have a direct line-of-sight to each target (i.e., not obscured by buildings) and satisfy other sensing constraints. For a gimbaled camera, the VV for a target located at ${\bm{g}}$ is the subset of the feasible airspace $F$ lies within direct line-of-sight to the target, and within a maximum range $d_{\rm max}$ relative to the target:
@@
For brevity, VVs (4) are denoted $\alpha({\bm{g}})$. The maximum range $d_{\rm max}$ is determined by the minimum pixel requirements for imaging a target of a known size for a desired resolution. VVs are computed using [4], resulting in a triangular mesh.
@@ 3.5 Problem Statement: Reference Orbit Design
An orbit may serve as a reference path for a UAV continuously encircling the POI. The reference orbit should be inside the POI’s VV as the POI travels through the environment over a time interval $[t_{0},t_{f}]$ where $t_{f}>t_{0}$. By following the VO, inside the VV, the UAV maintains a view of the target at all times. Let $O(\theta,t)~{}:~{}\mathbb{S}\times\mathbb{R}\to\mathbb{R}^{2}$ denote a time-varying orbit that is parameterized by the angle $\theta$ about the moving point $\bm{g}(t)=[g_{x}(t)~{}g_{y}(t)]^{\rm T}$ and the time $t$. For each point in time $t$, the orbit function $O(\theta,t)$ produces a closed non-self-intersecting curve in the $xy$ plane when sweeping out the angle $\theta\in\mathbb{S}$. The orbit is centered on ${\bm{g}}(t)$ and is located at a constant altitude $z=h_{\rm UAV}$. The curvature of the orbit in the inertial frame, denoted $\kappa(O(\theta,t),\theta,t)$ at any angle $\theta$ and time $t$, must remain less than the maximum curvature $\kappa_{\rm max}$ to be feasible for the system (1). The orbit must also remain in the POI’s VV, $\alpha(\bm{g}(t))$. While many types of orbit geometries are feasible, this work considers circular orbits. The time-varying orbit is defined as
@@
for $\theta\in\mathbb{S}$ where $R(t)$ is a time-varying radius. The standoff distance (i.e., the radius of the orbit) is maximized to reduce the curvature and observe the sides of the POI. The optimization problem is to specify $R(t)$ so that the orbit remains feasible for the Dubins vehicle model (1), guarantees the visibility of the POI, and maximizes the standoff distance:
@@
where $\bm{g}(t)$ is a known continuous path through the environment, $\bm{p}_{i}$ is the location of the $i$th (initial) node on the road network, and $\bm{p}_{j}$ is the final node. The first constraint (7) ensures the orbit is contained inside the VV. The curvature constraint (8) ensures that the UAV’s turn rate constraint (2) is satisfied on the orbit. The final constraints (9) and (10) specify the POI start and end positions on the road network $G$.
@@ 3.6 Problem Statement: Orbit Controller Design
The control objective is to design a controller $u_{\psi}(t,\bm{q})$ that steers the system (1) onto the orbit $O(\theta,t)$, $\lim_{t\rightarrow\infty}~{}[q_{x}(t),q_{y}(t)]^{\rm T}-O_{\rm min}(t,\bm{q}(t))\rightarrow\bm{0}$, where $O_{\rm min}(t,\bm{q}(t))\in\mathbb{R}^{2}$ is the closest point on the orbit to the configuration $\bm{q}(t)$, and the point $[q_{x}(t),q_{y}(t)]^{\rm T}$ is the inertial position component of the configuration $\bm{q}(t)$. In the case where the UAV is in the middle of the circular orbit, multiple points are equidistant from the internal position $[q_{x}(t),q_{y}(t)]^{\rm T}$ and the UAV is attracted to the point on the orbit in the direction of the UAV’s heading.
@@ 4 Moving View Polygons
First, this section shows the conditions under which a given orbit $R(t)$ (5) is feasible for the Dubins vehicle (1). Then, to better optimize the orbit $R(t)$, an adaptive discretization approach determines the locations $\mathcal{P}$ at which to compute the VVs, reducing changes in VV geometry. The VVs are then sliced along a constant altitude plane $z=h_{\rm UAV}$ with [24] to produce the VPs. Finally, an algorithm is presented to create the final time-varying orbit for the POI’s path while considering feasibility constraints. From this section forward, function parameters will be suppressed for brevity.
@@ 4.1 Feasibility of Reference Orbits
$\dot{\theta}=\frac{1}{R}\left(-\dot{\bm{g}}\cdot\bm{e}_{\theta}\pm\sqrt{v^{2}-(\dot{\bm{g}}\cdot\bm{e}_{r}+\dot{R})^{2}}\right)\;,$ (13)
where the “+” represents counter-clockwise (CCW) motion about $\bm{g}$ and the “-” represents clockwise (CW) motion.
@@
$\dot{\theta}=\frac{1}{R}\left(-\dot{\bm{g}}\cdot\bm{e}_{\theta}\pm\sqrt{(\dot{\bm{g}}\cdot\bm{e}_{r})^{2}-\left(v_{g}^{2}-v^{2}+2\dot{R}\dot{\bm{g}}\cdot\bm{e}_{r}+\dot{R}^{2}\right)}\right)\;.$ (19)
@@
Equation (13) is a time-varying differential equation valid over the interval $t\in[t_{i},t_{j}]$ for any initial condition $\theta(t_{i})=\theta_{i}$ and direction (CW or CCW) that defines a unique constant-speed trajectory remaining on the locus of orbits defined by $R$ and $\bm{g}$. The resulting path implicitly determines the vehicle heading on this orbit.
@@
Let “remain on the orbit $O$” be the condition wherein the angular velocity $\dot{\theta}$ about $\bm{g}$ is (13) and greater or equal to zero for CCW circulation and less than or equal to zero for CW circulation for all $\theta\in\mathbb{S}$.
@@
$\ddot{\bm{x}}_{\mathcal{I}}$ $=(\dot{a}-b\dot{\theta})\bm{e}_{r}+(\dot{b}+a\dot{\theta})\bm{e}_{\theta}\;,$ (22)
@@
There is a similar relation for $\dot{x}_{\mathcal{I},r}$ and $\dot{x}_{\mathcal{I},\theta}$ with $\dot{x}_{\mathcal{I},x}$ and $\dot{x}_{\mathcal{I},y}$, resulting in an expression for signed curvature,
$\kappa_{s}$ $=\frac{\ddot{x}_{\mathcal{I},\theta}\dot{x}_{\mathcal{I},r}-\ddot{x}_{\mathcal{I},r}\dot{x}_{\mathcal{I},\theta}}{\left(\dot{x}_{\mathcal{I},r}^{2}+\dot{x}_{\mathcal{I},\theta}^{2}\right)^{3/2}}\;.$ (24)
@@
If the system (1) remains on the orbit $O$ then the unsigned curvature of the path (3) is
@@
Lemma 2 applies to system (1), because system (1) has constant velocity; Therefore, the signed curvature of a trajectory is (31) and the unsigned curvature is
@@
If a trajectory $\bm{x}_{\mathcal{I}}$ has constant speed $v$ and bounded unsigned curvature $\kappa\leq\kappa_{\rm max}$ then the system (1) can follow the trajectory. The Dubins model can be written as (1) with $u_{\psi}=v\kappa_{s}$ Therefore, a trajectory is uniquely defined by an initial condition, $\bm{q}=[q_{x,0},q_{y,0},q_{\psi,0}]^{\rm T}$, and a curvature control $|\kappa_{s}|<\kappa_{\rm max}$.
@@
Let the distance from the orbit be $\tilde{r}=R-\sqrt{(g_{x}-q_{x})^{2}+(g_{y}-q_{y})^{2}}$. The first derivative of the distance is
@@
$\ddot{\tilde{r}}=\pm\dot{\theta}\dot{\bm{g}}\cdot\bm{e}_{\theta}v\left(v^{2}-\left(\dot{\bm{g}}\cdot\bm{e}_{\theta}+\dot{R}\right)^{2}\right)^{-1/2}\sin(q_{\psi}-\theta)-\dot{\theta}\dot{\bm{g}}\cdot\bm{e}_{\theta}\;.$ (46)
@@
$\ddot{\tilde{r}}$ =±$=\pm\dot{\theta}\dot{\bm{g}}\cdot\bm{e}_{\theta}v\left(v^{2}-\left(\dot{\bm{g}}\cdot\bm{e}_{\theta}+\dot{R}\right)^{2}\right)^{-1/2}\left(\pm\frac{1}{v}\sqrt{v^{2}-\left(\dot{\bm{g}}\cdot\bm{e}_{\theta}+\dot{R}\right)^{2}}\right)-\dot{\theta}\dot{\bm{g}}\cdot\bm{e}_{\theta}=0\;.$
@@
Suppose that the system (1) begins on an orbit $O$ with initial conditions (38)–(40). The system (1) remains on the orbit $O$ if and only if the curvature control is given by the left-hand side (LHS) of (31) and the inequality of (31) is satisfied.
@@
Part (ii): The initial condition ${\bm{q}}_{0}$ describes the polar angle $\theta_{0}$ and position on the orbit $O(\theta_{0},t_{0})$. From Lemma 4, if system (1) follows the curvature control (31) with the initial condition $\bm{q}_{0}$, then system (1) will remain on the orbit. ∎
@@
If orbit $O$ satisfies (31) then $\kappa$ is defined by the LHS of (31). For (31) to be real-valued it is required that $v^{2}-(\dot{\bm{g}}\cdot\bm{e}_{r}+\dot{R})^{2}\geq 0$. Moreover, this implies that $-v\leq\dot{R}+\dot{\bm{g}}\cdot\bm{e}_{r}\leq v$ and